# vHeat: Building Vision Models upon Heat Conduction

### v1

@@ Abstract
A fundamental problem in learning robust and expressive visual representations lies in efficiently estimating the spatial relationships of visual semantics throughout the entire image. In this study, we propose vHeat, a novel vision backbone model that simultaneously achieves both high computational efficiency and global receptive field. The essential idea, inspired by the physical principle of heat conduction, is to conceptualize image patches as heat sources and model the calculation of their correlations as the diffusion of thermal energy. This mechanism is incorporated into deep models through the newly proposed module, the Heat Conduction Operator (HCO), which is physically plausible and can be efficiently implemented using DCT and IDCT operations with a complexity of $\mathcal{O}(N^{1.5})$. Extensive experiments demonstrate that vHeat surpasses Vision Transformers (ViTs) across various vision tasks, while also providing higher inference speeds, reduced FLOPs, and lower GPU memory usage for high-resolution images. The code will be released at https://github.com/MzeroMiko/vHeat and https://openi.pcl.ac.cn/georgew/vHeat.
@@ 1 Introduction
When addressing these limitations, we draw inspiration from the field of heat conduction, where spatial locality is crucial for the transfer of thermal energy due to the collision of neighboring particles. Notably, analogies can be drawn between the principles of heat conduction and the propagation of visual semantics within the spatial domain, as adjacent image regions in a certain scale tend to contain related information or share similar characteristics. Leveraging these connections, we introduce vHeat, a physics-inspired vision backbone model that conceptualizes image patches as heat sources and models the calculation of their correlations as the diffusion of thermal energy.
To integrate the principle of heat conduction into deep networks, we first derive the general solution of heat conduction in 2D space and extend it to multiple dimensions, corresponding to various feature channels. Based on this general solution, we design the Heat Conduction Operator (HCO), which simulates the propagation of visual semantics across image patches along multiple dimensions. Notably, we demonstrate that HCO can be approximated through 2D DCT and IDCT operations, effectively reducing the computational complexity to $\mathcal{O}(N^{1.5})$. This improvement boosts both training and testing efficiency due to the high parallelizability of DCT and IDCT operations. Furthermore, as each element in the frequency domain obtained by DCT incorporates information from all patches in the image space, vHeat can establish long-range feature dependencies and achieve global receptive fields. To enhance the representation adaptability of vHeat, we propose learnable frequency value embeddings (FVEs) to characterize the frequency information and predict the thermal diffusivity of visual heat conduction.
@@
We develop a family of vHeat models (i.e., vHeat-Tiny/Small/Base), and extensive experiments are conducted to demonstrate their effectiveness in diverse visual tasks. Compared to benchmark vision backbones with various architectures (e.g., ConvNeXt [33], Swin [32], and Vim [67]), vHeat consistently achieves superior performance on image classification, object detection, and semantic segmentation across model scales. Specifically, vHeat-Base achieves a $83.9\%$ top-1 accuracy on ImageNet-1K, surpassing Swin by $0.4\%$, with a throughput exceeding that of Swin by a substantial margin over $40\%$ ($661$ vs. $458$). Besides, due to the $\mathcal{O}(N^{1.5})$ complexity of HCO, vHeat enjoys considerably lower computational cost compared to ViT-based models, demonstrating significantly reduced FLOPs and GPU memory requirements, and higher throughput as image resolution increases. In particular, when the input image resolution increases to $768\times 768$, vHeat-Base achieves a 3$\times$ throughput compared to Swin, with over 70% less GPU memory allocation and more than 30% fewer computational FLOPs.
@@ 2 Related Work
State Space Models and RNNs. State space models (SSMs) [20, 37, 55], which have the long-sequence modeling capacity with linear complexity, are also migrated from the natural language area (Mamba [19]). Visual SSMs were also designed by adapting the selective scan mechanism to 2-D images [67, 30]. Nevertheless, SSMs built upon the selective scan mechanism lack the advantages of high parallelism, which limit its potential.
Recent receptance weighted key value (RWKV) and RetNet models [38, 47] improved the parallelism while retaining the linear complexity. They combine the efficient parallelizable training of transformers with the efficient inference of RNNs, leveraging a linear attention mechanism and allowing formulation of the model as either a Transformer or an RNN, thus parallelizing computations during training and maintaining constant computational and memory complexity during inference. Despite the advantages, modeling a 2-D image as a sequence impairs interpretability.
@@ 3.1 Preliminaries: Physical Heat Conduction
where $k>0$ is the thermal diffusivity [3] measuring the rate of heat transfer in a material. By setting the initial condition $u(x,y,t)|_{t=0}$ to $f(x,y)$, the general solution of Eq. (1) can be derived by applying the Fourier Transform (FT, denoted as $\mathcal{F}$) to both sides of the equation, which gives
@@
$u(x,y,t)$ $=\mathcal{F}^{-1}(\widetilde{f}(\omega_{x},\omega_{y})e^{-k(\omega_{x}^{2}+\omega_{y}^{2})t})$ (7)
[…]
Figure 2: HCO and HCO layer. FVEs, FFN, LN, DWConv respectively denote frequency value embeddings, feed-forward network, layer normalization, and depth-wise convolution [^2] Please refer to Sec. E.3 in Appendix, where we demonstrate that while depth-wise convolution aids in feature extraction, the primary improvements are attributed to the proposed HCO..
@@ 3.2.1 Heat Conduction Operator (HCO)
Notably, due to the computational efficiency of $\mathbf{DCT_{2D}}$, the overall complexity of HCO is $\mathcal{O}(N^{1.5})$, where $N$ denotes the number of input image patches. Please refer to Sec. B in Appendix for the detailed implementation of HCO using $\mathbf{DCT_{2D}}$ and $\mathbf{IDCT_{2D}}$.
@@ 3.2.2 Adaptive Thermal Diffusivity
3.3 Discussion
[…]
Table 1: Performance comparison of image classification on ImageNet-1K. Test throughput values are measured with an A100 GPU, using the toolkit released by [61], following the protocol proposed in [32]. The batch size is set as 128, and the PyTorch version is 2.2. Please refer to Sec. D in Appendix for complete comparisons.
@@ 3.3 Discussion
ConvNeXt-S [33] 2242 50M 8.7G 684 83.1
ConvNeXt-B [33] 2242 89M 15.4G 436 83.8
Swin-T [32] 2242 28M 4.6G 1244 81.3
Swin-S [32] 2242 50M 8.7G 718 83.0
Swin-B [32] 2242 88M 15.4G 458 83.5
vHeat-T 2242 29M 4.6G 1514 82.2
vHeat-S 2242 50M 8.5G 945 83.6
vHeat-B 2242 87M 14.9G 661 83.9
Table 2: Left: Results of object detection and instance segmentation on COCO dataset. FLOPs are calculated with input size $1280\times 800$. $AP^{b}$ and $AP^{m}$ denote box AP and mask AP, respectively. The notation ‘1×1$\times$1 ×’ indicates models fine-tuned for 12 epochs, while ‘3×3$\times$3 ×MS’ denotes the utilization of multi-scale training for 36 epochs. Right: Results of semantic segmentation on ADE20K using UperNet [62]. FLOPs are calculated with the input size of $512\times 2048$.
@@
Swin-T 42.7 39.3 48M 267G
ConvNeXt-T 44.2 40.1 48M 262G
vHeat-T 45.1 41.0 53M 286G
Swin-S 44.8 40.9 69M 354G
ConvNeXt-S 45.4 41.8 70M 348G
vHeat-S 46.8 42.3 74M 377G
Swin-B 46.9 42.3 107M 496G
ConvNeXt-B 47.0 42.7 108M 486G
vHeat-B 47.7 43.0 115M 526G
@@
Swin-T 46.0 41.6 48M 267G
ConvNeXt-T 46.2 41.7 48M 262G
vHeat-T 47.2 42.4 53M 286G
Swin-S 48.2 43.2 69M 354G
ConvNeXt-S 47.9 42.9 70M 348G
vHeat-S 48.8 43.7 74M 377G
Crop size $\boldsymbol{512\times 512}$ on ADE20K
Backbone mIoU #param. FLOPs
ResNet-50 42.1 67M 953G
DeiT-S + MLN 43.8 58M 1217G
Swin-T 44.4 60M 945G
Vim-S 44.9 46M -
ConvNeXt-T 46.0 60M 939G
vHeat-T 46.9 65M 969G
ResNet-101 43.8 86M 1030G
DeiT-B + MLN 45.5 144M 2007G
Swin-S 47.6 81M 1039G
NAT-S 48.0 82M 1010G
ConvNeXt-S 48.7 82M 1027G
vHeat-S 49.0 86M 1062G
Swin-B 48.1 121M 1188G
NAT-B 48.5 123M 1137G
ConvNeXt-B 49.1 122M 1170G
vHeat-B 49.6 129M 1219G
4 Experiment
Experiments are performed to assess vHeat and compare it against popular CNN and ViT models. Visualization analysis is presented to gain deeper insights into the mechanism of vHeat. The evaluation spans three vision tasks including image classification on ImageNet-1K, object detection on COCO, and semantic segmentation on ADE20K. Please refer to Sec. C for experimental settings.
4.1 Image Classification
The image classification results are summarized in Table  1. With similar FLOPs, vHeat-T achieves a top-1 accuracy of $82.2\%$, outperforming DeiT-S by $2.4\%$, and Swin-T by $0.9\%$, respectively. Notably, the superiority of vHeat is also observed at both Small and Base scales. Specifically, vHeat-B achieves a top-1 accuracy of $83.9\%$, outperforming DeiT-B by $2.1\%$, and Swin-B by $0.4\%$, respectively.
In terms of computational efficiency, vHeat enjoys significantly higher inference speed across Tiny/Small/Base model scales compared to benchmark models. For instance, vHeat-T achieves a throughput of 1514 image/s, $87\%$ higher than Vim-S, $26\%$ higher than ConvNeXt-T, and $22\%$ higher than Swin-T, while maintaining a performance superiority, respectively.
4.2 Downstream Task
@@
Object Detection and Instance Segmentation. As a backbone network, vHeat is tested on the MS COCO 2017 dataset [29] for object detection and instance segmentation. We load classification pre-trained vHeat weights for downstream evaluation. Considering the input image size is different from the classification task, the shape of FVEs or $k$ should be aligned to the target image size on downstream tasks. Please refer to Sec. E.1 for ablation of interpolation for downstream tasks. The results for object detection are summarized in Table  2 (left), and vHeat enjoys superiority in box/mask Average Precision (APb and APm) in both of the training schedules ($12$ or $36$ epochs). For example, with a $12$-epoch fine-tuning schedule, vHeat-T/S/B models achieve object detection mAPs of $45.1\%/46.8\%/47.7\%$, outperforming Swin-T/S/B by $2.4\%/2.0\%/0.8\%$ mAP, and ConvNeXt-T/S/B by $0.9\%/1.4\%/0.7\%$ mAP, respectively. With the same configuration, vHeat-T/S/B achieve instance segmentation mAPs of $41.0\%/42.3\%/43.0\%$, outperforming Swin-T/S/B by $1.7\%/1.4\%/0.7\%$ mAP, and ConvNeXt-T/S/B by $0.9\%/0.5\%/0.3\%$ mAP, respectively. The advantages of vHeat persist under the $36$-epoch (3×3$\times$3 ×) fine-tuning schedule with multi-scale training. These results showcase vHeat’s potential to achieve promising performance in downstream tasks with dense prediction.
@@
Semantic Segmentation. The results on ADE20K are summarized in Table  2 (right), and vHeat consistently achieves superior performance. For example, vHeat-B respectively outperform Swin-B [32] and ConvNeXt-B [33] by 1.5% and 0.5% mIoU.
4.3 Visualization Analysis

### v2

# Building Vision Models upon Heat Conduction

@@ Abstract
Visual representation models leveraging attention mechanisms are challenged by significant computational overhead, particularly when pursuing large receptive fields. In this study, we aim to mitigate this challenge by introducing the Heat Conduction Operator (HCO) built upon the physical heat conduction principle. HCO conceptualizes image patches as heat sources and models their correlations through adaptive thermal energy diffusion, enabling robust visual representations. HCO enjoys a computational complexity of $O(N^{1.5})$, as it can be implemented using discrete cosine transformation (DCT) operations. HCO is plug-and-play, combining with deep learning backbones produces visual representation models (termed vHeat) with global receptive fields. Experiments across vision tasks demonstrate that, beyond the stronger performance, vHeat achieves up to a 3$\times$ throughput, 80% less GPU memory allocation, and 35% fewer computational FLOPs compared to the Swin-Transformer. Code is available at https://github.com/MzeroMiko/vHeat and https://openi.pcl.ac.cn/georgew/vHeat.
@@ 1 Introduction
When addressing these limitations, we draw inspiration from the field of heat conduction [70], where spatial locality is crucial for the transfer of thermal energy due to the collision of neighboring particles. Notably, analogies can be drawn between the principles of heat conduction and the propagation of visual semantics within the spatial domain, as adjacent image regions in a certain scale tend to contain related information or share similar characteristics. Leveraging these connections, we introduce vHeat, a physics-inspired vision representation model that conceptualizes image patches as heat sources and models the calculation of their correlations as the diffusion of thermal energy.
To integrate the principle of heat conduction into deep networks, we first derive the general solution of heat conduction in 2D space and extend it to multiple dimensions, corresponding to feature channels. Based on this general solution, we design the Heat Conduction Operator (HCO), which simulates the propagation of visual semantics across image patches along multiple dimensions. Notably, we demonstrate that HCO can be approximated through 2D (inverse) discrete cosine transformation (DCT/IDCT), effectively reducing the computational complexity to $\mathcal{O}(N^{1.5})$, Fig. 1. This improvement boosts both training and testing efficiency due to the high parallelizability of DCT and IDCT operations. Furthermore, as each element in the frequency domain obtained by DCT incorporates information from all patches in the image space, vHeat can establish long-range feature dependencies and achieve global receptive fields. To enhance the representation adaptability of vHeat, we propose learnable frequency value embeddings (FVEs) to characterize the frequency information and predict the thermal diffusivity of visual heat conduction.
We develop a family of vHeat models (i.e., vHeat-Tiny/Small/Base), and extensive experiments are conducted to demonstrate their effectiveness in diverse visual tasks. Compared to benchmark vision backbones with various architectures (e.g., ConvNeXt [39], Swin [37], and Vim [82]), vHeat consistently achieves superior performance on image classification, object detection, and semantic segmentation across model scales. Specifically, vHeat-Base achieves a $84.0\%$ top-1 accuracy on ImageNet-1K, surpassing Swin by $0.5\%$, with a throughput exceeding that of Swin by a substantial margin over $40\%$ ($661$ vs. $456$). To explore the generalization of vHeat, we’ve also validated its superiority on robustness evaluation benchmarks and low-level vision tasks. Besides, due to the $\mathcal{O}(N^{1.5})$ complexity of HCO, vHeat enjoys considerably lower computational cost compared to ViT-based models, demonstrating significantly reduced FLOPs and GPU memory requirements, and higher throughput as image resolution increases. In particular, when the input image resolution increases to $768\times 768$, vHeat-Base achieves a 3$\times$ throughput compared to Swin, with 80% less GPU memory allocation and 35% fewer computational FLOPs, as shown in Fig. LABEL:fig:cost.
@@ 2 Related Work
Figure 2: The network architecture of vHeat. Following the traditional principles of visual model design, we built vHeat with 4 HCO blocks, connected by downsampling layers in between.
@@
State Space Models and RNNs. State space models (SSMs) [22, 43, 65], which have the long-sequence modeling capacity with linear complexity, are also migrated from the natural language area (Mamba [21]). Visual SSMs were also designed by adapting the selective scan mechanism to 2-D images [82, 36]. Nevertheless, SSMs based on the selective scan mechanism suffer from limited parallelism, restricting their overall potential. Recent receptance weighted key value (RWKV) and RetNet models [44, 56] improved the parallelism while retaining the linear complexity. They combine the efficient parallelizable training of transformers with the efficient inference of RNNs, leveraging a linear attention mechanism and allowing formulation of the model as either a Transformer or an RNN, thus parallelizing computations during training and maintaining constant computational and memory complexity during inference. Despite the advantages, modeling a 2-D image as a sequence impairs interpretability.
@@ 3.1 Preliminaries: Physical Heat Conduction
where $k>0$ is the thermal diffusivity [5], measuring the rate of heat transfer in a material. By setting the initial condition $u(x,y,t)|_{t=0}$ to $f(x,y)$, the general solution of Eq. (1) can be derived by applying the Fourier Transform (FT, denoted as $\mathcal{F}$) to both sides of the equation, which gives
@@
$u(x,y,t)$ $=\mathcal{F}^{-1}(\widetilde{f}(\omega_{x},\omega_{y})e^{-k(\omega_{x}^{2}+\omega_{y}^{2})t}).$ (7)
Figure 3: HCO and HCO layer. FVEs, FFN, LN, and DWConv respectively denote frequency value embeddings, feed-forward network, layer normalization, and depth-wise convolution. Please refer to Sec. D.3 in the supplementary, where we demonstrate that while depth-wise convolution aids in feature extraction, the primary improvements are attributed to the proposed HCO.
@@ 3.2.1 Heat Conduction Operator (HCO)
Notably, due to the computational efficiency of $\mathbf{DCT_{2D}}$, the overall complexity of HCO is $\mathcal{O}(N^{1.5})$, where $N$ denotes the number of input image patches. Please refer to Sec. B in the supplementary for the detailed implementation of HCO using $\mathbf{DCT_{2D}}$ and $\mathbf{IDCT_{2D}}$.
@@ 3.2.2 Adaptive Thermal Diffusivity
Table 1: Performance comparison of image classification on ImageNet-1K. Test throughput values are measured with an A100 GPU, using the toolkit released by [71], following the protocol proposed in [37]. The batch size is set as 128, and the PyTorch version is 2.2.
@@
DCFormer-SW-T [33] 5122 28M 4.5G - 82.1
Vim-S [82] 2242 26M 5.3G 811 81.4
vHeat-T (Ours) 2242 29M 4.6G 1514 82.2
Swin-S [37] 2242 50M 8.7G 720 83.0
ConvNeXt-S [39] 2242 50M 8.7G 687 83.1
DCFormer-SW-S [33] 5122 50M 8.7G - 82.9
vHeat-S (Ours) 2242 50M 8.5G 945 83.6
Swin-B [37] 2242 88M 15.4G 456 83.5
ConvNeXt-B [39] 2242 89M 15.4G 439 83.8
RepLKNet-31B [16] 2242 79M 15.3G - 83.5
DCFormer-SW-B [33] 5122 88M 15.4G - 83.5
Vim-B [82] 2242 98M 19.0G 294 83.2
vHeat-B (Ours) 2242 68M 11.2G 661 84.0
3.3 Discussion
[…]
4 Experiment & Analysis
Experiments are performed to assess vHeat and compare it against popular CNN and ViT models. Visualization analysis is presented to gain deeper insights into the mechanism of vHeat. The evaluation spans image classification, object detection, semantic segmentation, out-of-distribution classification, and low-level vision tasks. Please refer to Sec. C in the supplementary for experimental settings.
4.1 Experimental Results
Image classification. The image classification results are summarized in Table  1. With similar FLOPs, vHeat-T achieves $82.2\%$ top-1 accuracy, outperforming Swin-T/Vim-S by $0.9\%/0.8\%$, respectively. Notably, the superiority of vHeat is also observed at both Small and Base scales. Specifically, vHeat-B achieves a top-1 accuracy of $84.0\%$ with only 11.2G FLOPs and 68M model parameters, outperforming Swin-B/Vim-B by $0.5\%/0.8\%$, respectively.
In terms of computational efficiency, vHeat enjoys significantly higher inference speed across Tiny/Small/Base model scales compared to benchmark models. For instance, vHeat-T achieves a throughput of 1514 images/s, $87\%$ higher than Vim-S, $26\%$ higher than ConvNeXt-T, and $22\%$ higher than Swin-T, while maintaining a performance superiority, respectively.
Table 2: Results of object detection and instance segmentation on COCO. FLOPs are calculated with input size $1280\times 800$. $AP^{b}$ and $AP^{m}$ denote box AP and mask AP, respectively. The notation ‘1×1$\times$1 ×’ indicates models fine-tuned for 12 epochs, while ‘3×3$\times$3 ×MS’ denotes the utilization of multi-scale training for 36 epochs.
@@ 4.1 Experimental Results
Swin-T 42.7 39.3 26.3 267G
ConvNeXt-T 44.2 40.1 29.3 262G
vHeat-T (Ours) 45.1 41.2 32.7 272G
Swin-S 44.8 40.9 19.7 359G
ConvNeXt-S 45.4 41.8 20.2 349G
vHeat-S (Ours) 46.8 42.3 25.9 348G
Swin-B 46.9 42.3 13.8 504G
ConvNeXt-B 47.0 42.7 14.1 486G
vHeat-B (Ours) 47.7 43.0 20.2 432G
@@
Swin-T 46.0 41.6 26.3 267G
ConvNeXt-T 46.2 41.7 29.3 262G
vHeat-T (Ours) 47.2 42.4 32.7 272G
Swin-S 48.2 43.2 19.7 359G
ConvNeXt-S 47.9 42.9 20.2 349G
vHeat-S (Ours) 48.8 43.7 25.9 348G
@@
Object Detection and Instance Segmentation. As a backbone network, vHeat is tested on the MS COCO 2017 dataset [35] for object detection and instance segmentation. We load classification pre-trained vHeat weights for downstream evaluation. Considering the input image size is different from the classification task, the shape of FVEs or $k$ should be aligned to the target image size on downstream tasks. Please refer to Sec. D.1 in the supplementary for ablation of interpolation for downstream tasks. The results for object detection are summarized in Table  2, and vHeat enjoys superiority in box/mask Average Precision (APb and APm) in both of the training schedules ($12$ or $36$ epochs). For example, with a $12$-epoch fine-tuning schedule, vHeat-T/S/B models achieve object detection mAPs of $45.1\%/46.8\%/47.7\%$, outperforming Swin-T/S/B by $2.4\%/2.0\%/0.8\%$ mAP, and ConvNeXt-T/S/B by $0.9\%/1.4\%/0.7\%$ mAP, respectively. With the same configuration, vHeat-T/S/B achieve instance segmentation mAPs of $41.2\%/42.3\%/43.0\%$, outperforming Swin-T/S/B and ConvNeXt-T/S/B. The advantages of vHeat persist under the $36$-epoch (3×3$\times$3 ×) fine-tuning schedule with multi-scale training. Besides, vHeat enjoys much higher inference speed (FPS) compared with Swin and ConvNeXt. For example, vHeat-B achieves 20.2 images/s, 46%/43% higher than Swin-B/ConvNeXt-B (13.8/14.1 images/s). These results highlight vHeat’s potential to deliver strong performance and efficiency in dense prediction downstream tasks.
Table 3: Results of semantic segmentation on ADE20K using UperNet [72]. FLOPs are calculated with the input size of $512\times 512$.
UperNet on ADE20K
Backbone mIoU FPS (images/s) FLOPs
Swin-T 44.4 31.8 237G
ConvNeXt-T 46.0 37.8 235G
ViL-S 46.3 - -
vHeat-T (Ours) 46.9 36.7 235G
Swin-S 47.6 22.1 261G
NAT-S 48.0 23.1 254G
ConvNeXt-S 48.7 27.7 257G
vHeat-S (Ours) 49.1 26.1 254G
Swin-B 48.1 19.2 299G
NAT-B 48.5 20.8 285G
ViL-B 48.8 - -
ConvNeXt-B 49.1 21.6 293G
vHeat-B (Ours) 49.6 23.6 293G
Table 4: Robust comparison of vHeat-B with Swin-B.
Model ObjectNet ImageNet-A
top-1 acc. (%) top-1 acc. (%)
Swin-B 25.4 36.0
ConvNeXt-B 26.1 36.5
vHeat-B (Ours) 26.7 36.8
@@
Semantic Segmentation. The results on ADE20K are summarized in Table  3, and vHeat consistently achieves superior performance over other baseline models across Tiny/Small/Base scales. For example, vHeat-B respectively outperform NAT-B [23] and ViL-B [2] by 1.1%/0.8% mIoU.
Robustness evaluation. To validate the robustness of vHeat, We evaluated vHeat-B on out-of-distribution classification datasets, including ObjectNet [3] and ImageNet-A [25]. We measure the Top-1 accuracy (%) for these two benchmarks, Table  4. It is evident that vHeat outperforms Swin and ConvNeXt consistently (better results are marked in bold). These experiments highlight vHeat’s robustness across out-of-distribution data, such as rotated objects, different view angles (ObjectNet), and natural adversarial examples (ImageNet-A).
Figure 5: Color image denoising visualization of vHeatIR and SwinIR after 15000 training iterations ($\sigma=15$). The input image is selected from McMaster [78].
Table 5: Quantitative comparison (average PSNR) on low-level vision tasks. : results are reproduced for a fair comparison.
Model Image Denoising JPEG Compression
(Set12/McMaster, Artifact Reduction
$\sigma=15$) (LIVE1, $q=40$)
DnCNN [76] 32.86/33.45 33.96
DRUNet [77] 33.25/35.40 34.58
SwinIR [34] 33.33/35.55 34.61
vHeatIR (Ours) 33.37/35.60 34.64
Low-level vision tasks. To further evaluate the generalization capability of our proposed vHeat model, we integrate the Heat Conduction Operator (HCO) by replacing the self-attention modules in the SwinIR [34] model, resulting in the vHeatIR architecture. We then conduct a series of experiments on several standard low-level vision tasks to assess the performance of vHeatIR. These tasks include grayscale and color image denoising on the Set12 [48] and McMaster [78] datasets, as well as JPEG compression artifact reduction on the LIVE1 [50] dataset. In these experiments, we use the same settings as those in SwinIR to ensure a fair comparison. The results, as summarized in Table 5, demonstrate that vHeatIR consistently outperforms the other baseline models. This improvement is largely attributed to the ability of HCO to operate efficiently in the frequency domain, which enhances the model’s performance in handling low-level image details. After training for 15,000 iterations, we visualize the denoising results on color images with noise level $\sigma=15$ in Fig. 5. As shown, vHeatIR produces noticeably cleaner images compared to SwinIR, indicating its superior ability to restore image quality. These results not only highlight the effectiveness of the proposed vHeat model but also validate its strong generalization capabilities for low-level vision tasks.
4.2 Analysis of Dynamic Locality